# Advanced safety filter based on SOS Control Barrier and Lyapunov Functions

### v2

@@ IV-E Alternating Algorithm
The solution of the SOS problem 39 can be found by iteratively alternate between searching over one set of decision variables while keeping the others fixed, see [11, 8]. In this process, either constraint margins are minimized or the volume of $\mathcal{X}_{s}$ is maximized. The volume is approximated by the summation of traces of $Q_{B_{i}}$ in $B_{i}(x)=Z_{B,i}(x)^{T}Q_{B,i}Z_{B,i}(x)$, similar to the approach in [7, 20]. Such an alternating algorithm requires a feasible initialization of the decision variables that are kept fixed in the first iteration. This is achieved through an initialization procedure that starts with an initially empty region of interest and expands it until it encompasses the allowable set $\mathcal{X}_{a}$. For more details, we refer the reader to Section IX.
@@ V Safety Filter Implementation
The objective of this section is to implement the advanced safety filter as specified in Problem 2. Given a linear and quadratic input constraints set $\mathcal{U}$, we propose the following realization of the Quadratically Constrained Quadratic Program (QCQP) in closed loop for every state $x\in\mathbb{R}^{n}$ along the trajectory:
$\begin{array}[]{llll}u_{s}(x):=&\underset{u\in\mathcal{U}}{\mbox{min}}&\|u_{n}(x)-u\|^{2}\\ &\mbox{s.t.}&C(x)u+b(x)\leq r(x),\\ \end{array}$ (42)
@@
$C(x)$ $:=\begin{bmatrix}\nabla V(x)^{T}G(x)\\ \nabla B_{1}(x)^{T}G(x)\\ \vdots\\ \nabla B_{t}(x)^{T}G(x)\\ \end{bmatrix}$ (43a)
$b(x)$ $:=\begin{bmatrix}\nabla V(x)^{T}f(x)+d(x)\\ \nabla B_{1}(x)^{T}f(x)\\ \vdots\\ \nabla B_{t}(x)^{T}f(x)\end{bmatrix}$ (43b)
encode the CLF and CBF conditions 15 and 17. The smooth functions $r(x)=\begin{bmatrix}r_{1}(x)&\ldots&r_{t}(x)\end{bmatrix}$ are state-dependent slack variables. Within the domain of CLF and CBF conditions 15 and 17, they must satisfy the upper bounds, given by:
@@
for all $i\in\mathcal{I}_{t}$ and $x\in\partial\mathcal{X}_{s}$. On the one hand, selecting a larger positive value for $r_{0}(x)$ within $\mathcal{X}_{n}$ (i.e. outside the domain of the CLF condition) increases the allowed dissipation rate violation in terms of larger values of $\dot{V}(x)=\nabla V(x)^{T}\dot{x}$. On the other hand, selecting a larger positive value for $r_{i}(x)$, $i\in\mathcal{I}_{t}$, within $\mathcal{X}_{s}$ (i.e. outside the domain of the CBF condition) increases the allowed safety violation in terms of larger positive values of $\dot{B}(x)=\nabla B(x)^{T}\dot{x}$. This renders the safety filter more passive within the interior of the safe set, and thus leaves more room to satisfy the CLF constraint. However, an overly large value for $r_{i}(x)$, $i\in\mathcal{I}_{t}$, can result in abrupt modifications of the control action when approaching the boundary of the safe set.
Remark 4
$\nabla V(x)^{T}(f(x)+G(x)u)\leq-\alpha(V(x)),$ (46a)
$\nabla B_{i}(x)^{T}(f(x)+G(x)u)\leq-\gamma_{i}(B_{i}(x))\quad i\in\mathcal{I}_{t}.$ (46b)
In this case, the class $\mathcal{K}$ functions $\alpha$ and $\gamma_{i}$ depend on the value of $V(x)$ and $B_{i}(x)$. However, in 43, we consider a broader class of the allowed dissipation rate violation $r_{0}(x)$ and allowed safety violation, $r_{i}(x)$, $i\in\mathcal{I}_{t}$, to ensure a smooth safety filter behavior, while satisfying further requirements presented below.
To ensure perfect tracking of the legacy controller $u_{n}(x)$ within $\mathcal{X}_{n}$ (cf. 9), we further lower bound $r(x)$ as:
@@
for all $x\in\mathcal{X}_{n}$. Indeed, condition 47 ensures that the cost of the QCQP in 42 can be minimized to zero within $\mathcal{X}_{n}$ with $u=u_{n}(x)$, making $u_{s}(x)$ coincide with $u_{n}(x)\in\mathcal{U}$. To ensure existence of $r(x)$ satisfying lower and upper bounds 44, 45 and 47, we need to show for $r_{0}(x)$ that $\nabla V(x)^{T}\left(f(x)+G(x)u_{n}(x)\right)+d(x)\leq 0$ on the intersection of $\mathcal{X}_{t}$ (cf. 44) and $\mathcal{X}_{n}$ (cf. 47). This is however directly ensured by forward invariance condition 19. Furthermore, we need to show for $r_{i}(x)$, $i\in\mathcal{I}_{t}$, that $\nabla B_{i}(x)^{T}\left(f(x)+G(x)u_{n}(x)\right)\leq 0$ on the intersection of $\partial\mathcal{X}_{s}$ (cf. 45) and $\mathcal{X}_{n}$ (cf. 47). This is however automatically satisfied as the domains do not overlap due to the containment condition 10.
Additionally, the selection of $r(x)$ must be lower-bounded to ensure the feasibility of the QCQP 42 within $\mathcal{X}_{t}$, as 47 already ensures feasibility within $\mathcal{X}_{n}$. Although the exact lower bound is unknown, it can be approximated using a feasible (but not necessarily optimal) solution $u_{\text{sos}}(x)$ from Section IV:
$C(x)u_{\text{sos}}(x)+b(x)\leq r(x),$ (48)
for all $x\in\mathcal{X}_{t}$. To guarantee the existence of such an $r(x)$, note that the lower bound 48 satisfies the upper bound conditions 44 and 45 as a result of the SOS constraints 25a and 25b.
@@
The Linear Independence Constraint Qualification (LICQ) is satisfied for the QCQP in 43 across all states $x$ and inputs $u$, that is, the gradients w.r.t. $u$ of the active inequality constraints in 43 are linearly independent. Equivalently, the rows of $C(x)$ associated with the active inequality constraints are linearly independent.
@@
By extending the QCQP in 43 with a slack variable $\delta\in\mathbb{R}^{t+1}$ as in [21]
$\begin{array}[]{llll}u_{s}(x):=&\underset{(u,\delta)\in\mathcal{U}\times\mathbb{R}^{t+1}}{\mbox{argmin}}&\|u_{n}(x)-u\|^{2}+\lambda\|\delta\|^{2}\\ &\mbox{s.t.}&C(x)^{T}u+b(x)\leq\delta,\\ \end{array}$
@@
The local Lipschitz-continuity not only preserves existence and uniqueness of solutions (cf. [15][Theorem 3.1]), but can also avoid the resulting feedback controller from chattering, as explained in [22]. The following theorem establishes local Lipschitz-continuity of the developed QCQP controller. Its proof can be found in the appendix.
@@
Under Assumption 3, given a smooth $B(x)$ and $V(x)$ as specified in Problem 1, a smooth legacy controller $u_{n}(x)$, and twice differentiable $r_{i}(x)$, $i\in\{0\}\cup\mathcal{I}_{t}$, lower bounded by 48, then $u_{s}(x)$ as defined in 43 is locally Lipschitz-continuous on the safe set $\mathcal{X}_{s}$.
We can find polynomial candidates $r(x)$ based on given CBF $B(x)$ and CLF $V(x)$ by solving an additional SOS problem that encodes the upper and lower bounds 44, 45, 47 and 48.
@@ VI Power Converter Case Study
Figure 5: The schematic is presented in the (dq) rotating frame, compromising a DC-link, 3-phase bridge, and an ac-line connected to an infinite bus $v_{g}$. Assuming the average-value model for the 3-phase bridge, the input $\bar{u}=\begin{bmatrix}m_{d}&m_{q}\end{bmatrix}^{T}$ determines the converter voltage $v_{c}$ as a function of the dc-link voltage $v_{dc}$. Likewise, the dc-current flowing into the capacitor $c_{dc}$ is given by the input $\bar{u}$ and the ac-line current $i=\begin{bmatrix}i_{d}&i_{q}\end{bmatrix}^{T}$. The states consisting of the dc-link voltage $v_{dc}$ and the line-current $i$ are measured.
@@
where $l=0.01$ is the inductance in p.u. consisting of transformer and line inductance, $c_{dc}=0.02$ is the dc-link capacitance in p.u., and $g_{dc}:=1/r_{dc}=0.001$ is the admittance of the DC-link capacitor in p.u. The measured states are denoted by $\bar{x}:=\begin{bmatrix}v_{dc}&i_{d}&i_{q}\end{bmatrix}^{T}$, and the input is given by $\bar{u}:=\begin{bmatrix}m_{d}&m_{q}\end{bmatrix}^{T}$, describing the modulation indices. The objective is to stabilize both the DC-link voltage $v_{dc}$ to $1$, and the quadrature current $i_{q}$ to its reference $i_{q,ref}$, which is assumed to be constant. As a consequence, to meet the steady-state specifications, the direct current $i_{d}$ is stabilized to $-g_{dc}$. The allowable set of states must adhere to the following inequality constraints:
@@
Furthermore, the quadratic input constraint $\mathcal{U}$ is given by
$m_{d}^{2}+m_{q}^{2}\leq(1.2)^{2}.$ (51)
@@
We introduce the error coordinates for the states $x=\bar{x}-x^{*}=\begin{bmatrix}\tilde{v}_{dc}&\tilde{i}_{d}&\tilde{i}_{q}\end{bmatrix}^{T}$ and the input variables $u=\bar{u}-u^{*}=\begin{bmatrix}\tilde{m}_{d}&\tilde{m}_{q}\end{bmatrix}^{T}$ such that $\dot{x}=0$ at $x=0$ and $u=0$. The system represented in error coordinates is described by system 2 with
@@
In the following subsection, we first consider a single quadrature current reference set to $i_{q,ref}=0$. Then, we introduce a quadrature current reference as a new state to the dynamical system, with a zero time-derivative, i.e., ${\dot{i}}_{q,ref}=0$. For each scenario, we find the CBFs $B_{i}(x)$ and CLF $V(x)$ (cf. Problem 1), and perform simulations that illustrate the additional guarantees provided by the advanced safety filter $u_{s}(x)$ compared with a standard safety filter implementation. We use CVXOPT Python library to solve the SDPs of the alternating algorithm, and to implement the QCQP in 42. This library implements an optimization problem solver based on second-order and positive semidefinite cones. The simulations are performed using the ODE solver from scipy with a sampling time of $1$ $\mu s$. The full code used for simulations and optimization is publicly available online [^1] https://github.com/MichaelSchneeberger/advanced-safety-filter/.
Figure 6: The allowable set $\mathcal{X}_{a}$, the safe set $\mathcal{X}_{s}$, and the nominal region $\mathcal{X}_{n}$ projected to the $\tilde{v}_{dc}$ and $\tilde{i}_{d}$ coordinate are shown for $\tilde{i}_{q}=i_{q,ref}=0$. In addition, the zero-level sets of the polynomials $B_{1}(x)$ and $B_{2}(x)$ are shown in grey. Three scenarios are shown, where (a) the input action is subjected to no constraints, i.e. $u(x(t))\in\mathbb{R}^{m}$, (b) the input action is subjected to a quadratic constraint given by $u(x(t))\in\mathcal{U}:=\{u\mid u^{T}u\leq 1.3\}$, and (c) $i_{q,ref}$ is integrated as a new state of the dynamical system using the same input constraints. The corresponding vector field $\dot{x}=f(x)+G(x)u_{sos}(x)$ (projected to $(\tilde{v}_{dc},\tilde{i}_{d})$ coordinates) is shown in blue, where $u_{sos}(x)=p(x)/s(x)$ is used to ensure compatibility between the CBF and CLF.
@@ VI-A Zero quadrature current reference tracking
In this subsection, we assume a zero quadrature current reference, i.e., $i_{q,ref}=0$. The allowable set $\mathcal{X}_{a}$, as expressed in 12, encoding the state constraints 50, are achieved through the following polynomials:
$w_{1}(x)$ $=((\tilde{v}_{dc}+0.3)/0.5)^{2}+(\tilde{i}_{d}/20)^{2}+(\tilde{i}_{q}/20)^{2}-1$
$w_{2}(x)$ $=(\tilde{v}_{dc}/20)^{2}+({\tilde{i}_{d}}/1.3)^{2}+({\tilde{i}_{q}}/1.3)^{2}-1.$
Figures 6 (a) and (b) illustrates the safe set $\mathcal{X}_{s}$ and the nominal set $\mathcal{X}_{n}$ computed by the algorithm both without and with considering inputs constraints. Table I summarizes both the number of variables associated with the respective SDPs and the number of constraints, derived from the summation of triangular elements within the positive semidefinite matrices. The initialization procedure completed in 10 iterations, where each iteration took around 10 seconds. The main algorithm completed in 20 iterations, where each iteration took around 14 seconds. The degree of the polynomials are given in Table II.
@@
The proposed safety filter introduces additional guarantees, which make it more stringent and, as noted in the introduction, more suitable in already established industrial setups. However, due these system requirements, it now represents a new category that prioritizes the preservation of the legacy control, enhancing composability, preventing a direct performance comparison between conventional safety filters and our novel setup. Nonetheless, side-by-side plots in Figure 7 are included that illustrate the additional guarantees compared to a basic safety filter in selected scenarios. The advanced safety filter $u_{s}(x)$ is based on the previously computed CLF $V(x)$ and CBFs $B_{1}(x)$ and $B_{2}(x)$, and is compared to a basic safety filter $u_{\text{SF}}(x)$ using the same CBFs:
$\begin{array}[]{lll}\underset{u\in\mathcal{U}}{\mbox{argmin}}&\|u_{n}(x)-u\|^{2}\\ \mbox{s.t.}&\nabla B_{1}(x)(f(x)+G(x)u)+\alpha\left(B_{1}(x)\right)\leq 0\\ &\nabla B_{2}(x)(f(x)+G(x)u)+\alpha\left(B_{2}(x)\right)\leq 0,\end{array}$ (57)
where $\alpha(B_{i}(x)):=10B_{i}(x)$ is selected.
Figure 7: (Convergence to the nominal region $\mathcal{X}_{n}$): While the advanced safety filter $u_{s}(x)$ essentially overlaps the legacy control $u_{n}$, the basic safety filter $u_{\text{SF}}(x)$ suboptimally perturbs the legacy control action, resulting in a much slower convergence.
Moreover, a side-by-side plots in Figure 8 illustrate the effect on including quadratic input constraints into the design of the safety filter. We consider the integration of quadratic input constraints into the SOS constraint as a contribution distinct from the discussion on the safety filter’s role in preserving the legacy control action. However, for consistency, the simulations were performed using the previous setup, which including an advanced safety filter and the basic safety filter in 57. Since the input constraints are incorporated into the design of the CLF and CBFs, the safety filter $u_{s}(x)$ successfully stays within the state constraints for all $t\geq 0$. This is not true for a basic safety filter $u_{\text{SF}}(x)$, which does not consider the input constraints in its design. In this case, the norm of the input action $\|u_{\text{SF}}(x)\|$ exceeds the input constraint limit of $1.3$ p.u. in the beginning of the simulation (cf. Figure 8). To produce a valid input, the input action $u_{\text{SF}}(x)$ is projected to the set of feasible input set $\mathcal{U}$. As a result, the trajectory deviates from its intended path, leading to a violation of the dc-link voltage constraint.
Figure 8: (Forward invariance of the safe set $\mathcal{X}_{s}$): While the advanced safety filter $u_{s}(x)$, which considers the quadratic input constraints into the design, successfully stays within the state constraints, whereas the legacy controller $u_{n}$ and basic safety filter $u_{\text{SF}}(x)$ violate the state constraints.

### v3

@@ IV-E Alternating Algorithm
The solution of the SOS problem 40 can be found by iteratively alternate between searching over one set of decision variables while keeping the others fixed, see [19, 18]. In this process, either constraint margins are minimized or the volume of $\mathcal{X}_{s}$ is maximized. The volume is approximated by the summation of traces of the gram matrix $Q_{B_{i}}$ in $B_{i}(x)=Z_{B,i}(x)^{\top}Q_{B,i}Z_{B,i}(x)$, similar to the approach in [23, 12]. Such an alternating algorithm requires a feasible initialization of the decision variables that are kept fixed in the first iteration. This is achieved through an initialization procedure that starts with an initially empty region of interest and expands it until it encompasses the allowable set $\mathcal{X}_{a}$. For more details, we refer the reader to the supplementary material of this paper.
@@ V Safety Filter Implementation
Given a linear input constraints set $\mathcal{U}$, we propose the following realization of the QP in closed loop for every state $x\in\mathbb{R}^{n}$ along the trajectory:
[…]
@@
$C(x):=\left[\begin{smallmatrix}\nabla V(x)^{\top}G(x)\\ \nabla B_{1}(x)^{\top}G(x)\\ \vdots\\ \nabla B_{n_{\!B}}(x)^{\top}G(x)\\ \end{smallmatrix}\right],\,b(x):=\left[\begin{smallmatrix}\nabla V(x)^{\top}f(x)+d(x)\\ \nabla B_{1}(x)^{\top}f(x)\\ \vdots\\ \nabla B_{n_{\!B}}(x)^{\top}f(x)\end{smallmatrix}\right]$ (44a)
encode the CLF and CBF conditions 16 and 18. The smooth functions $r(x)=[\begin{matrix}r_{1}(x)&\ldots&r_{n_{\!B}}(x)\end{matrix}]$ represent state-dependent slack variables that specify the allowed dissipation rate and safety violation. These slack-variables can be found using an SOS optimization, subject to upper and lower bounds that enforce the specifications outlined in Problem 2.
Remark 4
$\nabla V(x)^{\top}(f(x)+G(x)u)\leq-\alpha(V(x)),$ (45a)
$\nabla B_{i}(x)^{\top}(f(x)+G(x)u)\leq-\gamma_{i}(B_{i}(x))\quad i\in\mathcal{I}_{n_{\!B}}.$ (45b)
In this case, the class $\mathcal{K}$ functions $\alpha$ and $\gamma_{i}$ depend on the value of $V(x)$ and $B_{i}(x)$. However, in 44, we consider a broader class of functions $r_{0}(x)$ and $r_{i}(x)$, $i\in\mathcal{I}_{n_{\!B}}$, to ensure the preservation of the legacy control behavior during nominal conditions.
V-A Upper bounds on $r(x)$
Within the domain of CLF and CBF conditions 16 and 18, the slack-variables must satisfy the upper bounds, given by
@@ V-A Upper bounds on $r(x)$
for all $i\in\mathcal{I}_{n_{\!B}}$ and $x\in\partial\mathcal{X}_{s}$.
V-B Lower bounds on $r(x)$
To ensure tracking of the legacy controller $u_{n}(x)$ within $\mathcal{X}_{n}$ (cf. 10), we impose the following lower bound on $r(x)$:
@@ V-B Lower bounds on $r(x)$
for all $x\in\mathcal{X}_{n}$. This condition guarantees that the cost in the QP in 43 can be minimized to zero within $\mathcal{X}_{n}$ by choosing $u=u_{n}(x)$, ensuring that the safety filter $u_{s}(x)$ coincides with the legacy control action $u_{n}(x)\in\mathcal{U}$. The existence of $r_{0}(x)$ satisfying 46 and 48 is ensured through 20, while the existence of $r_{i}(x)$, $i\in\mathcal{I}_{n_{\!B}}$, satisfying 47 and 48 is ensured if $\mathcal{X}_{n}\subseteq\text{int}(\mathcal{X}_{s}$).
Furthermore, the selection of $r(x)$ must be lower-bounded to ensure the feasibility of the QP 43 within $\mathcal{X}_{s}$. Although the exact lower bound is unknown, it can not be directly obtain as it involves an existence quantifier $\exists u\in\mathcal{U}$ for every state $x$. However, it can be approximated using a feasible (but not necessarily optimal) solution $u_{\text{sos}}(x)$ from Section IV:
$C(x)u_{\text{sos}}(x)+b(x)\leq r(x)$ (49)
for all $x\in\mathcal{X}_{t}$. The existence of $r(x)$ satisfying 46, 47 and 49 is guaranteed through the SOS constraints 26a and 26b. Typical lower and upper bound selections for $r_{0}(x)$ are illustrated in the supplementary material of this paper.
V-C Resulting SOS Problem
The polynomial candidate $r(x)$ satisfying 46, 47, 48 and 49 can be obtained by solving the SOS problem:
[…]
where $B_{i}$, $V$, $p$, $s$ are taken from solving the SOS problem in 40, and the gram matrix $Q_{r,i}$ is defined by $r_{i}(x)=Z(x)^{\top}Q_{r,i}Z(x)$. The cost function in 58 favors choices of $r(x)$ that yield smooth intervention of the safety filter as the system approaches the state constraints.
V-D Lipschitz-continuity
@@
The Linear Independence Constraint Qualification (LICQ) is satisfied for the QP in 44 across all states $x$ and inputs $u$, that is, the gradients w.r.t. $u$ of the active inequality constraints in 44 are linearly independent. Equivalently, the rows of $C(x)$ associated with the active inequality constraints are linearly independent.
@@
By extending the QP in 44 with a slack variable $\delta\in\mathbb{R}^{t+1}$ as in [3]
[…]
@@
The local Lipschitz-continuity not only preserves existence and uniqueness of solutions (cf. [10][Theorem 3.1]), but can also avoid the resulting feedback controller from chattering, as explained in [15]. The following theorem establishes local Lipschitz-continuity of the developed QP controller. Its proof is given in Section VIII-E.
@@
Under Assumption 3, given a smooth $B(x)$ and $V(x)$ as specified in Problem 1, a smooth legacy controller $u_{n}(x)$, and twice differentiable $r_{i}(x)$, $i\in\{0\}\cup\mathcal{I}_{n_{\!B}}$, lower bounded by 49, then $u_{s}(x)$ as defined in 44 is locally Lipschitz-continuous on the safe set $\mathcal{X}_{s}$.
@@ VI Power Converter Case Study
Figure 5: The schematic is presented in the (dq) rotating frame, compromising a DC-link, 3-phase bridge, and an ac-line connected to an infinite bus $v_{g}$. Assuming the average-value model for the 3-phase bridge, the input $\bar{u}=\begin{bmatrix}m_{d}&m_{q}\end{bmatrix}^{\top}$ determines the converter voltage $v_{c}$ as a function of the dc-link voltage $v_{dc}$. Likewise, the dc-current flowing into the capacitor $c_{dc}$ is given by the input $\bar{u}$ and the ac-line current $i=\begin{bmatrix}i_{d}&i_{q}\end{bmatrix}^{\top}$. The states consisting of the dc-link voltage $v_{dc}$ and the line-current $i$ are measured.
@@
where $l=0.01$ is the inductance in p.u. consisting of transformer and line inductance, $c_{dc}=0.02$ is the dc-link capacitance in p.u., and $g_{dc}:=1/r_{dc}=0.001$ is the admittance of the DC-link capacitor in p.u. The measured states are denoted by $\bar{x}:=[\begin{matrix}v_{dc}&i_{d}&i_{q}\end{matrix}]^{\top}$, and the input is given by $\bar{u}:=[\begin{matrix}m_{d}&m_{q}\end{matrix}]^{\top}$, describing the modulation indices. The objective is to stabilize both the DC-link voltage $v_{dc}$ to 11, and the quadrature current $i_{q}$ to its reference $i_{q,ref}$, which is assumed to be constant. As a consequence, to meet the steady-state specifications, the direct current $i_{d}$ is stabilized to $g_{dc}$. The allowable set of states must adhere to the following inequality constraints:
@@
We introduce the error coordinates for the states $x=\bar{x}-x^{*}=[\begin{matrix}\tilde{v}_{dc}&\tilde{i}_{d}&\tilde{i}_{q}\end{matrix}]^{\top}$ and the input variables $u=\bar{u}-u^{*}=[\begin{matrix}\tilde{m}_{d}&\tilde{m}_{q}\end{matrix}]^{\top}$ such that $\dot{x}=0$ at $x=0$ and $u=0$. The system represented in error coordinates is described by system $\dot{x}=f(x)+G(x)u$ with
@@
In the following subsection, we first consider a single quadrature current reference set to $i_{q,ref}=0$. Then, we introduce a non-zero quadrature current reference as a new state to the dynamical system. The full code used for the optimization is available online [^1] https://github.com/MichaelSchneeberger/advanced-safety-filter/.
Figure 6: The allowable set $\mathcal{X}_{a}$, the safe set $\mathcal{X}_{s}$, and the nominal region $\mathcal{X}_{n}$ projected to the $\tilde{v}_{dc}$ and $\tilde{i}_{d}$ coordinate are shown for $\tilde{i}_{q}=i_{q,ref}=0$. In addition, the zero-level sets of the polynomials $B_{1}(x)$ and $B_{2}(x)$ are shown in gray. Three scenarios are shown, where (a) the input action is subjected to no constraints, i.e. $u(x(t))\in\mathbb{R}^{m}$, (b) the input action is subjected to a quadratic constraint given by $u(x(t))\in\mathcal{U}:=\{u\mid u^{\top}u\leq 1.2\}$, and (c) $i_{q,ref}$ is integrated as a new state of the dynamical system using the same input constraints. The corresponding vector field $\dot{x}=f(x)+G(x)u_{sos}(x)$ (projected to $(\tilde{v}_{dc},\tilde{i}_{d})$ coordinates) is shown in blue, where $u_{sos}(x)=p(x)/s(x)$ is used to ensure compatibility between the CBF and CLF.
@@ VI-A Zero quadrature current reference tracking
In this subsection, we assume a zero quadrature current reference, i.e., $i_{q,ref}=0$. The allowable set $\mathcal{X}_{a}$, as expressed in 13, encoding the state constraints 60, are achieved through the following polynomials:
$w_{1}(x)$ $=((\tilde{v}_{dc}+0.3)/0.5)^{2}-1$
$w_{2}(x)$ $=({\tilde{i}_{d}}/1.3)^{2}+({\tilde{i}_{q}}/1.3)^{2}-1.$
Figures 6 (a) and (b) illustrates the safe set $\mathcal{X}_{s}$ and the nominal region $\mathcal{X}_{n}$ computed by the algorithm both without and with considering inputs constraints. Table I summarizes both the number of variables associated with the respective SDPs and the number of constraints, derived from the summation of triangular elements within the positive semidefinite matrices. The initialization procedure completed in 10 iterations, where each iteration took around 10 seconds. The main algorithm completed in 20 iterations, where each iteration took around 14 seconds. The degree of the polynomials are given in Table II. Simulation results comparing the advanced safety filter with a basic safety filter are provided in the supplementary material.